OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (aakashsudhakar):

x1 = 2; y1 = -5; m = -3. y - y1 = m(x - x1) y + 5 = -3(x - 2) y + 5 = -3x + 6 y = -3x + 6 - 5 = -3x + 1 y = -3x + 1 It's the last option.